Bob is 7 years older than ALi.4 years time ,Bob will be twice as old as Ali.how old is Ali now.
Answer by Cromlix(4381)

Hi there,
Bob is 7 years older than ALi
B = 7 + A
4 years time ,Bob will be twice as old as Ali
B + 4 = 2(A + 4)
Replace B with (7 + A)
7 + A + 4 = 2(A + 4)
7 + A + 4 = 2A + 8
Collect like terms
A - 2A = 8 - 4 - 7
- A = - 3
A = 3
Ali is 3 years old
Bob is 10 years old.
Hope this helps:-)